# BrainFLORA: Uncovering Brain Concept Representation via Multimodal Neural Embeddings

### v1

@@ Abstract.
Understanding how the brain represents visual information is a fundamental challenge in neuroscience and artificial intelligence. While AI-driven decoding of neural data has provided insights into the human visual system, integrating multimodal neuroimaging signals—such as EEG, MEG, and fMRI—remains a critical hurdle due to their inherent spatiotemporal misalignment. Current approaches often analyze these modalities in isolation, limiting a holistic view of neural representation. In this study, we introduce BrainFLORA, a unified framework for integrating cross-modal neuroimaging data to construct a shared neural representation. Our approach leverages multimodal large language models (MLLMs) augmented with modality-specific adapters and task decoders, achieving state-of-the-art performance in joint-subject visual retrieval task and has the potential to extend multitasking. Combining neuroimaging analysis methods, we further reveal how visual concept representations align across neural modalities and with real-world object perception. We demonstrate that the brain’s structured visual concept representations exhibit an implicit mapping to physical-world stimuli, bridging neuroscience and machine learning from different modalities of neural imaging. Beyond methodological advancements, BrainFLORA offers novel implications for cognitive neuroscience and brain-computer interfaces (BCIs). Our code is available at https://github.com/ncclab-sustech/BrainFLORA.
Neuroscience, Neural Coding, Concept-Selective, Visual Decoding, BCI, Multi-modality, Generative Model
*Co-first author: Dongyang Li and Haoyang Qin (lidy2023@mail.sustech.edu.cn, 12211832@mail.sustech.edu.cn).
‡Corresponding author: Quanying Liu (liuqy@sustech.edu.cn).
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; Oct 27–31, 2025; Dublin, Irelandccs: Human-centered computing HCI design and evaluation methodsccs: Computing methodologies Artificial intelligenceccs: Cognitive science
@@ 1. Introduction
One core challenge in achieving advanced performance lies in establishing cross-modality representations for diverse neural data. Such data exhibit significant complexity and variability, influenced by factors such as individual subjects, acquisition devices, visual content, and spatial-temporal dynamics. Recent works such as NeuroBind (yang2024neurobind, ), which deploys distinct encoders for each type of neural data. Additionally, one-encoder architectures like UMBRAE (xia2024umbrae, ) have aimed to learn a uniform representation across subjects in fMRI for brain decoding tasks. While these efforts demonstrate progress in improving performance for various downstream tasks, they remain constrained by the inherent limitations of effective modal alignment. This challenge highlights the critical need for more cohesive and innovative frameworks that not only integrate multiple modalities but also facilitate each unique modal characteristic to enhance understanding of brain representations related to different neural signals.
@@ 3. Method
Formulation. Let $T$ represent the length of neural data, $C$ the number of channels, and $N$ the total number of data samples. Our objective is to derive neural embeddings $Z=\mathcal{E}(X)\in\mathbb{R}^{N\times F}$ from the brain activity data $X\in\mathbb{R}^{N\times C\times T}$ and when $X$ is used as fMRI modality, the dimension $T$ is not included, where $\mathcal{E}$ is the unified projector and $F$ is the projection dimension of the embeddings. Concurrently, we use the CLIP model to extract image embeddings $\hat{Z}\in\mathbb{R}^{N\times F}$ from images $I$. Our goal is to effectively align the multimodal representation with the image representation, as illustrated in Fig. 2.
@@ 3.1. Model Architecture
Using a small network $g$, we use the soft router computes a score vector for each token, with a length equal to the number of experts $K$. Finally, the Softmax function is applied to obtain the final routing vector, yielding the final routing vector. It can be formulated as:
@@ 3.2. Training Objective
Where $p_{i}=\mathcal{E}(x_{i}),t_{i}=\text{CLIP}_{\text{Image}}(y_{i})$, $\tau$ is a temperature hyperparameter, and $x_{i}$ and $y_{i}$ represent the $i-th$ neural data sample and image respectively. For generative tasks, including captioning and reconstruction, we incorporate Mean Squared Error (MSE) loss to ensure consistent model training for regression-based tasks. As a result, the overall loss function of BrainFLORA is formulated as a composite of these individual loss components. Inspired by MindEye (scotti2024reconstructing, ), the optimization of model parameters $\theta$ is then governed by the following objective function:
(13) $\min_{\theta}\mathcal{L}=\mathcal{L}_{\text{BiMixCo\textbar SoftCLIP}}+\alpha\cdot\mathcal{L}_{\text{MSE}}+\beta\cdot\mathcal{L}_{\text{prior}}$
@@
(14) $\mathcal{L_{\text{lowlevel}}}=\|D(f)-D(\hat{f})\|+\|f-\hat{f}\|+\lambda_{P}\mathcal{L}_{P}(D(\hat{f}))$
@@ 4.1. Dataset and Implementation
Table 1. Multimodal Retrieval Performance. We present the retrieval performance of various methods on the THINGS-EEG2, THINGS-MEG and THINGS-fMRI datasets. Retrieval refers to the task of correctly identifying paired neural latent representations from the 200 or 100 class test set clip embeddings. Performance is evaluated across joint subject retrieval tasks, and we report results for several retrieval configurations: 2-way, 4-way, 10-way, as well as Top-1 and Top-5 accuracy for the 200-way retrieval task (100 for fMRI). BrainFLORA is compared with previous SOTA methods, including NICE (song2023decoding, ), EEGNetV4 (lawhern2018eegnet, ), B.D. (benchetrit2024brain, ), MindEyeV2 (scottimindeye2, ), UMBRAE (xia2024umbrae, ), ATM-S (li2024visual, ), CongnitionCapturer (zhang2024cognitioncapturer, ), MindBridge (wang2024mindbridge, ), WAVE (wang2024decoding, ) and MB2C (wei2024mb2c, ). In-subject results refers to Appendix.
[…]
Table 2. Multimodal Visual Reconstruction. We quantitatively evaluated reconstruction quality across EEG, MEG, and fMRI modalities, reporting averaged results from all subjects. Unlike previous works that perform training and inference within individual subjects, BrainFLORA operates in a cross-subject, cross-modality setting for both training and reconstruction evaluation. Additional evaluation details can be found in Appendix Sec.B.
@@ 4.5. In-depth Analysis
To assess the fidelity of the resulting embedding, we conducted an evaluation by exhaustively gathering concept embedding for a group of all objects not included in the BrainFLORA’s training data. By computing their similarity, we compared this measured similarity matrix with the one predicted by the model’s embedding in Fig. 4. Significant positive correlations were observed between the model-predicted and measured Representational Similarity Matrices (RSMs) (correlation coefficients: 0.49 for EEG, 0.55 for MEG, and 0.4 for fMRI). These results validate that the 42-dimensional embeddings we analyzed closely reflect the similarity space underlying the corresponding visual concept embedding.
@@ 5. Discussion and Conclusion
These advances not only push the boundaries of neural decoding but also shed light on the intricate interactions between diverse neural data modalities. Moving forward, two critical directions emerge: 1) leveraging larger-scale neural datasets—both paired and unpaired—to train robust multimodal foundation models, and 2) pioneering grand unified architectures that integrate representation learning, modality alignment, and fusion mechanisms for generalizable neural decoding.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (62472206), Shenzhen Science and Technology Innovation Committee (RCYX20231211090405003,
KJZD20230923115221044), GuangDong Basic and Applied Basic Research Foundation (2025A1515011645 to ZC.L.), Shenzhen Doctoral Startup Project (RCBS20231211090748082 to XK.S.),

### v2

@@ Abstract.
Understanding how the brain represents visual information is a fundamental challenge in neuroscience and artificial intelligence. While AI-driven decoding of neural data has provided insights into the human visual system, integrating multimodal neuroimaging signals—such as EEG, MEG, and fMRI—remains a critical hurdle due to their inherent spatiotemporal misalignment. Current approaches often analyze these modalities in isolation, limiting a holistic view of neural representation. In this study, we introduce BrainFLORA, a unified framework for integrating cross-modal neuroimaging data to construct a shared neural representation. Our approach leverages multimodal large language models (MLLMs) augmented with modality-specific adapters and task decoders, achieving state-of-the-art performance in joint-subject visual retrieval task and has the potential to extend multitasking. Combining neuroimaging analysis methods, we further reveal how visual concept representations align across neural modalities and with real-world object perception. We demonstrate a fundamental alignment between the representational structures of visual objects as understood by AI and those found in the human brain, showing a shared framework for comprehending the physical world. Beyond methodological advancements, BrainFLORA offers novel implications for cognitive neuroscience and brain-computer interfaces (BCIs). Our code is available at https://github.com/ncclab-sustech/BrainFLORA.
Neuroscience, Neural Coding, Concept-Selective, Visual Decoding, BCI, Multimodality, Generative Model
Co-first authors: Dongyang Li and Haoyang Qin.
Corresponding authors: Chen Wei and Quanying Liu.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3754996ccs: Human-centered computing HCI design and evaluation methodsccs: Computing methodologies Artificial intelligenceccs: Cognitive science
@@ 1. Introduction
One core challenge in achieving advanced performance lies in establishing cross-modal representations for diverse neural data. Such data exhibit significant complexity and variability, influenced by factors such as individual subjects, acquisition devices, visual content, and spatial-temporal dynamics. Recent works such as NeuroBind (yang2024neurobind, ), which deploys distinct encoders for each type of neural data. Additionally, one-encoder architectures like UMBRAE (xia2024umbrae, ) have aimed to learn a uniform representation across subjects in fMRI for brain decoding tasks. While these efforts demonstrate progress in improving performance for various downstream tasks, they remain constrained by the inherent limitations of effective modal alignment. This challenge highlights the critical need for more cohesive and innovative frameworks that not only integrate multiple modalities but also facilitate each unique modal characteristic to enhance understanding of brain representations related to different neural signals.
@@ 3. Method
Formulation. Let $T$ represent the length of neural data, $C$ the number of channels, and $N$ the total number of data samples. Our objective is to derive neural embeddings $Z=\mathcal{E}(X)\in\mathbb{R}^{N\times F}$ from the brain activity data $X\in\mathbb{R}^{N\times C\times T}$ and when $X$ is used as fMRI modality, the dimension $T$ is not included, where $\mathcal{E}$ is the unified projector and $F$ is the projection dimension of the embeddings. Concurrently, we use the CLIP model to extract image embeddings $\hat{Z}\in\mathbb{R}^{N\times F}$ from images $I$. Our goal is to effectively align multimodal neural representations with image representations, as illustrated in Fig. 2.
@@ 3.1. Model Architecture
Using a small network $g$, we use the soft router computes a score vector for each token, with a length equal to the number of experts $K$. Finally, a Softmax function is applied to obtain the final routing vector, yielding the final routing vector. It can be formulated as:
@@ 3.2. Training Objective
Where $p_{i}=\mathcal{E}(x_{i}),t_{i}=\text{CLIP}_{\text{Image}}(y_{i})$, $\tau$ is a temperature hyperparameter, and $x_{i}$ and $y_{i}$ represent the ithi-th neural data sample and image respectively. For generative tasks, including captioning and reconstruction, we incorporate Mean Squared Error (MSE) loss to ensure consistent model training for regression-based tasks. As a result, the overall loss function of BrainFLORA is formulated as a composite of these individual loss components. Inspired by MindEye (scotti2024reconstructing, ), the optimization of model parameters $\theta$ is then governed by the following objective function:
[…]
@@
(14) $)\mathcal{L_{\text{lowlevel}}}=\|D(f)-D(\hat{f})\|+\|f-\hat{f}\|+\lambda_{P}\mathcal{L}_{P}(D(\hat{f}$))
@@ 4.1. Dataset and Implementation
Table 1. Multimodal Retrieval Performance. We present the retrieval performance of various methods on the THINGS-EEG2, THINGS-MEG and THINGS-fMRI datasets. Retrieval refers to the task of correctly identifying paired neural latent representations from the 200 or 100 class test set clip embeddings. Performance is evaluated across joint subject retrieval tasks, and we report results for several retrieval configurations: 2-way, 4-way, 10-way, as well as Top-1 and Top-5 accuracy for the 200-way retrieval task (100 for fMRI). BrainFLORA is compared with previous SOTA methods, including NICE (song2023decoding, ), EEGNetV4 (lawhern2018eegnet, ), B.D. (benchetrit2024brain, ), MindEyeV2 (scottimindeye2, ), UMBRAE (xia2024umbrae, ), ATM-S (li2024visual, ), CognitionCapturer (zhang2024cognitioncapturer, ), MindBridge (wang2024mindbridge, ), WAVE (wang2024decoding, ) and MB2C (wei2024mb2c, ). In-subject results refers to Appendix.
Model THINGS-EEG2 THINGS-MEG THINGS-fMRI 2-way 4-way 10-way 200-way 200-way 2-way 4-way 10-way 200-way 200-way 2-way 4-way 10-way 100-way 100-way Top-1 Top-1 Top-1 Top-1 Top-5 Top-1 Top-1 Top-1 Top-1 Top-5 Top-1 Top-1 Top-1 Top-1 Top-5 NICE 90.10 77.15 57.15 13.80 35.80 60.62 39.62 20.87 2.50 7.88 87.00 74.67 55.00 22.00 51.00 EEGNetV4 88.15 77.05 59.10 12.15 35.85 71.13 48.63 29.38 3.88 11.88 86.00 74.67 55.33 18.00 44.67 B.D. 90.05 76.85 55.85 12.45 34.15 62.12 35.50 17.63 1.88 6.50 72.67 58.33 38.00 6.67 26.00 MindEyeV2 90.25 79.85 64.30 17.35 43.65 64.25 40.25 20.37 2.13 8.38 92.00 80.33 64.33 21.33 60.33 UMBRAE 66.50 41.95 21.85 2.90 10.00 61.75 35.50 16.63 2.00 6.50 85.67 70.67 50.67 16.67 42.67 ATM-S 93.35 83.70 68.60 19.70 48.60 77.00 59.25 38.88 6.25 20.12 92.67 81.67 65.33 24.00 55.33 CognitionCapturer 92.90 84.00 67.85 20.50 50.70 67.88 47.63 23.57 2.38 10.37 86.67 73.00 55.33 20.67 48.00 MindBridge 89.10 75.35 57.45 15.00 39.15 64.88 40.12 20.13 1.87 7.37 88.33 73.33 55.67 16.67 54.33 WAVE 89.40 75.85 58.80 15.00 39.05 67.50 44.75 24.00 3.00 11.38 88.67 71.67 52.33 20.33 45.33 MB2C 85.20 70.25 49.65 10.20 29.85 61.75 33.25 17.00 1.25 5.50 84.67 69.00 49.67 18.67 42.33 BrainFLORA-uni 95.55 86.90 73.45 25.35 57.30 81.75 64.50 46.62 8.00 24.38 91.33 79.00 63.00 26.33 57.33 BrainFLORA-multi 94.05 87.30 73.15 25.05 56.35 80.50 61.88 39.75 6.88 23.38 92.33 84.67 70.67 28.33 63.33
Table 2. Multimodal Visual Reconstruction. We quantitatively evaluated reconstruction quality across EEG, MEG, and fMRI modalities, reporting averaged results from all subjects. Unlike previous works that perform training and inference within individual subjects, BrainFLORA operates in a cross-subject, cross-modal setting for both training and reconstruction evaluation. Additional evaluation details can be found in Appendix Sec.B.
@@ 4.5. In-depth Analysis
To assess the fidelity of the resulting embedding, we conducted an evaluation by exhaustively gathering concept embedding for a group of all objects not included in the BrainFLORA’s training data. By computing their similarity, we compared this measured similarity matrix with the one predicted by the model’s embedding in Fig. 4. Significant positive correlations were observed between the model-predicted and measured Representational Similarity Matrices (RSMs) (correlation coefficients: 0.50 for EEG, 0.56 for MEG, and 0.40 for fMRI). These results validate that the 42-dimensional embeddings we analyzed closely reflect the similarity space underlying the corresponding visual concept embedding.
@@ 5. Discussion and Conclusion
These advances not only push the boundaries of neural decoding but also shed light on the intricate interactions between diverse neural data modalities. Moving forward, two critical directions emerge: (1) leveraging larger-scale neural datasets—both paired and unpaired—to train robust multimodal foundation models, and (2) pioneering grand unified architectures that integrate representation learning, modality alignment, and fusion mechanisms for generalizable neural decoding.